# Class Distribution Shifts in Zero-Shot Learning: Learning Robust Representations

### v3

@@ 4 Proposed Approach
Motivated by the analysis of the parametric model we propose a new approach to tackle class distribution shifts in zero-shot learning. Our approach revolves around two key ideas: (i) during training, different mixtures of the attribute $A$ can be achieved through sampling of small subsets of the classes, and (ii) penalizing for differences in performance across these environments is likely to increase robustness to the class mixture encountered at testing.
@@ 4.1 Synthetic Environments
Standard ERM training involves sampling pairs of data points $(z_{i},z_{j})$ uniformly at random from all $N_{c}$ classes available during training. However, as discussed in §3, this is prone to overfitting to the training attribute mixture. Since the identity of the attribute is unknown, weighted sampling (and similar approaches) cannot be used to create environments with different attribute mixtures.
Yet, our goal is to design artificial environments with diverse compositions of the (unknown) attribute of interest. To do so we leverage the variability in small samples: while class subsets of similar size to $N_{c}$ maintain attribute distributions similar to the overall training set, smaller subsets with $k\ll N_{c}$ classes are likely to exhibit distinct attribute distributions. Therefore, we propose creating multiple environments, composed of examples from a few sampled classes.
This results in a hierarchical sampling scheme for the data pairs: first, sample a subset of $k$ classes, $S=\{c_{1},\dots,c_{k}\}$. Then, for each $c\in S$ sample $2r$ pairs of data points as follows: $r$ pairs from within the class $c$, $\{z_{i};c_{i}=c\}$, uniformly at random (positive pairs); and $r$ negative pairs, where one point is uniformly at random from $c$, and the other from all other data points in $S$, $\{z_{i};c_{i}\neq c,c_{i}\in S\}$. [^1] Here for simplicity we create balanced environments, but different proportions of positive examples can be considered instead.
Across multiple class subsets $S=\{S_{1},\dots,S_{n}\}$, this hierarchical sampling results in diverse mixtures of any unknown attribute (see Figure 3). In particular, in some of the class subsets, classes from the overall minority type form the majority within the environment.
Figure 3: Illustration of the proposed hierarchical sampling. Top: $N_{c}=6$ classes, with 2 minority-type classes D, F (in purple). Middle: synthetic environments formed by sampling small ($k=3$) class sets – in $\nicefrac{{1}}{{5}}$ of the environments, minority-type classes become the majority, constituting $\nicefrac{{2}}{{3}}$ of the classes. Bottom: sampling $r=1$ positive and $r=1$ negative pairs for each class in the environment.
@@ 4.2 Environment Balancing Algorithm for Class Distribution Shifts
To achieve this, inspired by performance balancing methods in OOD (see §2) we optimize a penalized objective:
@@
where $R\left(S_{1},\dots,S_{n}\right)$ is any balancing term across the constructed synthetic environments.
Note that computing $R\left(S_{1},\dots,S_{n}\right)$ often involves evaluating some value on each environment separately. For a general balancing term, we denote the value in the $l$-th environment as $f(S_{l})$, and accordingly express $R\left(S_{1},\dots,S_{n}\right)=\mathring{f}\left(f(S_{1}),\dots,f(S_{n})\right)$, where $\mathring{f}$ represents the corresponding aggregation function. Our combined approach [^2] For notation simplicity we assume the unpenalized training loss is applied to pairs of data points $(x_{ij},y_{ij})=((z_{i},z_{j}),\mathbbm{1}_{c_{i}=c_{j}})$, but it can be easily adapted for any tuple size (e.g., triplets)., which entails balancing performance over synthetic environments of class subsamples, is outlined in Algorithm 1.
@@
        Sample $k$ classes from $C^{*}$ without replacement: $S_{l}^{(t)}=\{c_{l}^{(1)},\dots,c_{l}^{(k)}\}$
@@
     Compute $\!R^{(t)}\!\coloneqq\!R\left(S_{1}^{(t)},\dots,S_{n}^{(t)}\right)\!=\!\mathring{f}\left(f(S_{1}^{(t)}),\dots,f(S_{n}^{(t)})\right)$
@@ 4.3 Balancing Performance Instead of Loss
In numerous OOD penalties, such as IRM and VarREx, $f$ is the loss in each environment, which is distance-based in deep metric learning algorithms. This presents a challenge in zero-shot verification, where sampled tuples often include numerous easy negative examples, leading to quick performance plateaus, despite considerable variations in distances themselves. Strategies like selecting the most difficult tuples [18] were proposed to address this issue. However, these methods have been observed to generate noisy gradients and loss values [34].
We therefore propose to directly balance performance instead of relying on the losses in the training environments. Denote the set of negative pairs within a synthetic environment by $D_{l}^{0}=\{x_{ij}=(z_{i},z_{j}):c_{i},c_{j}\in S_{l},y_{ij}=0\}$, and the set of positive pairs by $D_{l}^{1}=\{x_{ij}=(z_{i},z_{j}):c_{i},c_{j}\in S_{l},y_{ij}=1\}$. An unbiased estimator of the AUC on a given synthetic environment $S_{l}$ is given by
@@
for $x_{ij}\in D_{l}^{1}$ and $x_{uv}\in D_{l}^{0}$. Since this estimator is non-differentiable, and therefore cannot be used in gradient-descent based optimization, we use an approximation named soft-AUC [7]
@@ 4.4 How Many Environments are Needed?
The proposed hierarchical sampling allows for the construction of many synthetic environments with various attribute mixtures, influenced by the number of classes in each environment. As shown in the subsequent analysis, this ensures that, with high probability, classes with rare attribute values will appear together in the same environment. This, in turn, facilitates learning to separate negative pairs involving rare class types.
At each training iteration, we consider $n$ class subsets (environments) of size $k$. Our goal is to achieve robustness to all attribute values $a$ that are associated with at least $\rho_{\text{min}}\in(0,1)$ of the training classes. Note that $\rho_{\text{min}}$ is specified by the practitioner without knowledge of the true attribute that may cause the shift, or its true abundance $\rho$ in the training set.
We compute the number of synthetic environments $n$, such that with high probability of $(1-\alpha)$, $S_{1},\dots,S_{n}$ will include at least one subset with at least two classes associated with $a$ (otherwise none of the subsets would contain negative pairs with the attribute $a$). Denote the probability of a given subset to not contain any class associated with $a$ by $\phi_{0}=\nicefrac{{{\left\lceil(1-\rho_{\text{min}})N_{c}\right\rceil\choose k}}}{{{N_{c}\choose k}}}$, and the probability of a given subset to contain exactly one such class by $\phi_{1}=\nicefrac{{\rho_{\text{min}}N_{c}{\left\lceil(1-\rho_{\text{min}})N_{c}\right\rceil\choose k-1}}}{{{N_{c}\choose k}}}$. This corresponds to $n$ that is typically much smaller than (${N_{c}\choose k}$, given by
Figure 4: Average AUC over 10 simulation repetitions for $\rho=0.9$. Solid lines: distribution-shift. Dashed lines: in-distribution. Our method improves robustness for shifts, without compromising training distribution results.
@@ 5 Empirical Results
Our method enhances standard training with two components: hierarchical sub-sampling and a balancing term for synthetic environments. To our knowledge, this is the first work addressing OOD generalization for class distribution and zero-shot learning. We therefore benchmark our algorithm against the ERM baseline (uniform random sampling with an unpenalized score) and optimization of an unpenalized score over hierarchical samples, providing an ablation study. Additionally, we tested standard regularization techniques such as dropout and the $L_{2}$ norm, which did not yield notable improvements in the distribution shift scenario, and therefore are not shown.
For comprehensive comparisons, in addition to the proposed VarAUC penalty, we evaluate variants of our algorithm with IRM, CLOvE, and VarREx penalties as well. While we show that VarAUC consistently outperforms other penalties, the crucial improvement lies in its performance compared to the ERM baseline: application of existing OOD penalties is enabled by the construction of synthetic environments in our framework. As discussed in Appendix C, this construction facilitates the formulation of previously untreated class distribution shifts in zero-shot learning within the OOD setting.
Across all experiments, we trained the network with contrastive loss (Equation 2) and normalized cosine distance: $d_{g}(z_{1},z_{2})=\frac{1}{2}\left(1-\frac{g(z_{1})\cdot g(z_{2})}{\left\|g(z_{1})\right\|\left\|g(z_{2})\right\|}\right)$. The exact setups are detailed below and additional details appear in Appendix F.
@@ 5.1 Simulations: Revisiting the Parametric Model
Here we revisit the parametric model presented in §3. To increase the complexity of the problem we add dimensions where classes from both types are not well separated. That is, $\Sigma_{a}$ includes additional dimensions set to zero.
Setup We used 68 subsets in each training iteration, each consisting of two classes. This corresponds to choosing $\rho_{\text{min}}=0.1$ (desired sensitivity, regardless of the true unknown parameter $\rho\in\{0.05,0.1,0.3\}$) with a low $\alpha$ value of 0.5, resulting in the construction of fewer environments according to Equation 4.4. For each class, we sampled $102r=$ 10 pairs of data points. The representation was defined as $g(z)=wz$ for $w\in\mathbb{R}^{d\times p}$  [^3] A linear representation is chosen to facilitate an analysis of the learned representation space.. Here we focus on the case of $p=16$, $\nu_{z}=\nu_{0}=1$, $\nu^{-}=0.1$, $\nu_{+}=2$, $d_{0}=5$, $d_{1}=d_{2}=10$. Additional representation sizes $p$, noise ratios $\frac{\nu^{+}}{\nu^{-}}$ and varying proportions of positive and negative examples are shown in Appendix D.1.
@@
Results In Figure 5 we examine the learned representation. The analysis reveals that ERM prioritizes dimensions 5-15, providing good separation for $a_{1}$, the dominant type in training,
Figure 5: Avg. feature importance for $\rho=0.9$, 10 repetitions. Our VarAUC method emphasizes shared features (blocks 1 and 3) most, while de-prioritizing majority features (block 2). All methods assign low weight to noise features (block 4).
@@
In Figure 4, the learning progress is depicted for $\rho=0.9$, and similar analysis for $\rho=0.95$ and $\rho=0.7$ can be found in Appendix D . Performance on the same distribution as the training data is similar between ERM and our algorithm, suggesting that applying our algorithm does not negatively impact performance when no distribution shift occurs. However, under distribution shift our algorithm achieves much better results. The VarREx penalty achieves high AUC values more quickly than the VarAUC penalty, but VarAUC attains higher accuracy overall. IRM shows noisier convergence since it is applied directly on the gradients, which have been shown to be noisy in contrastive learning due to high variance in data-pair samples [34]. Means and standard deviations are reported in Appendix D.1, as well as results for additional data dimensions, positive proportions and variance ratios.
@@ 5.2 Experiments on Real Data
Experiment 1 - Face Recognition We used the CelebA dataset [30] that contains 202,599 images of 10,177 celebrities. We filtered out people for who less than 3 images were present. Following Vinyals et al. [51] we implemented $g$ as a convolutional neural network which has 4 modules with $3\times 3$ convolutions and 64 filters, followed by batch normalization, a ReLU activation, $2\times 2$ max-pooling and a fully connected layer of size 32. We used the attribute blond hair for the class distribution shift: for training we sampled mostly non-blond people ($95\%$) while in test most people ($95\%$) were blond. Training iteration had 150 synthetic environments of 2 classes and $202r=$ 20 data points per class.
Experiment 2 - Species Recognition We used the ETHEC dataset [11] which includes 47,978 butterfly images from 6 families and 561 species (example images are provided in Appendix D). We filtered out species with less than 5 images and focused on images of butterflies from the Lycaenidae and Nymphalidae families. We included in the training set $10\%$ of species from the Nymphalidae family, while at test time $90\%$ of the species were from the Nymphalidae family. For each class we sampled $202r=$ 20 pairs. We trained the models on 200 synthetic environments at a time, each of two classes. We implemented $g$ as a fully connected neural network with layers of sizes 128, 64, 32 and 16, and ReLU activations between them.
Figure 6: Average percentage changes across 10 repetitions compared to ERM are shown for the ETHEC (top) and CelebA (bottom) datasets. Error bars represent $\pm$ one std-dev.
Experimental Results Figure 6 shows that all versions of our algorithm achieve some improvement over ERM, where best results are achieved with VarAUC penalty. Exact means and standard deviations are reported in Table 3 in the Appendix. One-sided paired t-tests show that the improvement over ERM achieved by our algorithm with the VarAUC penalty is statistically significant, with p-values of $<0.04$ on both datasets. p-values for other penalties are reported in Table 4. All p-values were adjusted with FDR [4] correction.
In Appendix D we provide additional analysis confirming that the main improvement of our algorithm over the ERM baseline stems from improved performance on negative minority pairs.

### v4

@@ 4 Proposed Approach
Motivated by our analysis of the parametric model, we propose a new approach for tackling class distribution shifts in zero-shot learning. Our approach revolves around two key ideas: (i) during training, different mixtures of the attribute $A$ can be produced by sampling small subsets of the classes, forming artificial environments, and (ii) penalizing for differences in performance across these environments is likely to increase robustness to the class mixture encountered at test time.
@@ 4.1 Synthetic Environments
Standard ERM training involves sampling pairs of data points $(z_{i},z_{j})$ uniformly at random from all $N_{c}$ classes available during training. However, as discussed in §3, this is prone to overfitting to the attribute distribution of the training data. Since the identity of the attribute is unknown, weighted sampling (and similar approaches) cannot be used to create environments with different attribute mixtures.
Yet, our goal is to design artificial environments with diverse compositions of the (unknown) attribute of interest. To do so, we leverage the variability in small samples: while class subsets of similar size to $N_{c}$ maintain attribute mixtures similar to the overall training set, smaller subsets with $k\ll N_{c}$ classes are likely to exhibit distinct attribute mixtures. Therefore, we propose creating multiple environments, composed of examples from few sampled classes.
This results in a hierarchical sampling scheme for the data pairs: first, sample a subset of $k$ classes, $S=\{c_{1},\dots,c_{k}\}$. Then, for each $c\in S$ sample $2r$ pairs of data points as follows: $r$ pairs from within the class $c$, $\{z_{i};c_{i}=c\}$, uniformly at random (positive pairs); and $r$ negative pairs, where one point is sampled uniformly at random from $c$, and the other from all other data points in $S$, $\{z_{i};c_{i}\neq c,c_{i}\in S\}$. [^1] Here, for simplicity we create balanced environments, but different proportions of positive examples can be considered instead.
Across multiple class subsets $S=\{S_{1},\dots,S_{n}\}$, this hierarchical sampling results in diverse mixtures of any unknown attribute (see Figure 3). In particular, in some of the class subsets, classes from the overall minority type constitute the majority in the environment.
Figure 3: Illustration of the proposed hierarchical sampling. Top: $N_{c}=6$ classes, with 2 minority-type classes D, F (in purple). Middle: synthetic environments formed by sampling small ($k=3$) class subsets; in $\nicefrac{{1}}{{5}}$ of the environments, minority-type classes become the majority constituting $\nicefrac{{2}}{{3}}$ of the classes. Bottom: sampling $r=1$ positive and $r=1$ negative pairs for each class in the environment.
@@ 4.2 Environment Balancing Algorithm for Class Distribution Shifts
To achieve this, inspired by OOD performance balancing methods (see §2), we optimize a penalized objective:
@@
where $R\left(S_{1},\dots,S_{n}\right)$ is any balancing term between the constructed synthetic environments.
Note that computing $R\left(S_{1},\dots,S_{n}\right)$ often involves evaluating some value on each environment separately. For a general balancing term, we denote the value in the $l$-th environment as $f(S_{l})$ and accordingly express $R\left(S_{1},\dots,S_{n}\right)=\mathring{f}\left(f(S_{1}),\dots,f(S_{n})\right)$, where $\mathring{f}$ represents the corresponding aggregation function. Our approach [^2] For notation simplicity we assume that the unpenalized training loss is applied to pairs of data points $(x_{ij},y_{ij})=((z_{i},z_{j}),\mathbbm{1}_{c_{i}=c_{j}})$, but it can easily be adapted for any tuple size (e.g., triplets). for balancing performance across synthetic environments of class subsets, is outlined in Algorithm 1.
@@
        Sample $k$ classes from $C^{*}$ without replacement: $S_{l}^{(t)}=\{c_{l}^{(1)},\dots,c_{l}^{(k)}\}$.
@@
     Compute $\!R^{(t)}\!\coloneqq\!R\left(S_{1}^{(t)},\dots,S_{n}^{(t)}\right)\!=\!\mathring{f}\left(f(S_{1}^{(t)}),\dots,f(S_{n}^{(t)})\right)$.
@@ 4.3 Balancing Performance Instead of Loss
In multiple OOD penalties (e.g., IRM and VarREx), $f$ represents the loss in each environment, which, in deep metric learning algorithms, is based on distance. This presents a challenge in zero-shot verification, where sampled tuples often include numerous easy negative examples, leading to performance plateau early in the learning process, although the distances themselves still exhibit considerable variations. Strategies like selecting the most difficult tuples [18] were proposed to address this issue, however these methods have been found to generate noisy gradients and loss values [34].
We therefore propose to balance performance directly instead of relying on the losses in the training environments. Denote the set of negative pairs in a synthetic environment by $D_{l}^{0}=\{x_{ij}=(z_{i},z_{j}):c_{i},c_{j}\in S_{l},y_{ij}=0\}$ and the set of positive pairs by $D_{l}^{1}=\{x_{ij}=(z_{i},z_{j}):c_{i},c_{j}\in S_{l},y_{ij}=1\}$. An unbiased estimator of the AUC on a given synthetic environment $S_{l}$ is given by
@@
for $x_{ij}\in D_{l}^{1}$ and $x_{uv}\in D_{l}^{0}$. Since this estimator is non-differentiable and therefore cannot be used in gradient-descent-based optimization, we use soft-AUC as an approximation [7]
@@ 4.4 How Many Environments Are Needed?
The proposed hierarchical sampling scheme allows for the construction of many synthetic environments with various attribute mixtures, influenced by the number of classes in each environment. As shown in the analysis below, this ensures that with high probability there will be at least one environment with a pair of minority type classes, thereby supporting learning to separate negative pairs within the minority type.
In each training iteration, we consider $n$ class subsets (environments) of size $k$. Our goal is to achieve robustness to all attribute values $a$ that are associated with at least $\rho_{\text{min}}\in(0,1)$ of the training classes. Note that $\rho_{\text{min}}$ is specified by the practitioner without knowledge of the true attribute that may cause the shift or its true prevalence $\rho$ in the training set.
We compute the number of synthetic environments $n$, such that with high probability of $(1-\alpha)$, $S_{1},\dots,S_{n}$ will include at least one subset with at least two classes associated with $a$ (otherwise none of the subsets would contain negative pairs with the attribute $a$). Denote the probability of a given subset not to contain any class associated with $a$ by $\phi_{0}=\nicefrac{{{\left\lceil(1-\rho_{\text{min}})N_{c}\right\rceil\choose k}}}{{{N_{c}\choose k}}}$ and the probability of a given subset to contain exactly one such class by $\phi_{1}=\nicefrac{{\rho_{\text{min}}N_{c}{\left\lceil(1-\rho_{\text{min}})N_{c}\right\rceil\choose k-1}}}{{{N_{c}\choose k}}}$. Therefore, the required number of environments needed to ensure that at least two minority-type classes appear together in the same environment is
@@
Note that that $n$ is typically much smaller than (${N_{c}\choose k}$.
@@ 5 Empirical Results
Our method enhances standard training with two components: a hierarchical sampling scheme and a balancing term for synthetic environments. To the best of our knowledge, this is the first work addressing OOD generalization for class distribution in zero-shot learning. We therefore benchmark our algorithm against the ERM baseline (uniform random sampling with an unpenalized score) and a hierarchical sampling baseline (hierarchical sampling with unpenalized score).
Figure 4: Average AUC over 10 simulation repetitions for majority attribute proportion $\rho=0.9$ in training (and 0.1 in test). Solid lines: distribution-shift. Dashed lines: in-distribution. Our method improves robustness for shifts, without compromising training distribution results.
Additionally, we tested standard regularization techniques including dropout and the $L_{2}$ norm, which did not yield notable improvements in the distribution shift scenario, and are therefore not shown.
To ensure a comprehensive comparison, in addition to the proposed VarAUC penalty, we evaluate variants of our algorithm in which the IRM, CLOvE, and VarREx penalties are used instead. While we show that VarAUC consistently outperforms other penalties, the crucial improvement lies in its performance compared to the ERM baseline: application of existing OOD penalties is enabled by the construction of synthetic environments in our algorithm. As discussed in Appendix C, this construction facilitates the formulation of class distribution shifts in zero-shot learning within the OOD setting.
In all of the experiments performed, we trained the network with contrastive loss (Equation 2) and the normalized cosine distance: $d_{g}(z_{1},z_{2})=\frac{1}{2}\left(1-\frac{g(z_{1})\cdot g(z_{2})}{\left\|g(z_{1})\right\|\left\|g(z_{2})\right\|}\right)$. The specific setups are detailed below (additional details can be found in Appendix F), and code to reproduce our results is available at https://github.com/YuliSl/Zero_Shot_Robust_Representations .
@@ 5.1 Simulations: Revisiting the Parametric Model
We now revisit the parametric model presented in §3. To increase the complexity of the problem, we add dimensions where classes from both types are not well separated. That is, $\Sigma_{a}$ includes additional dimensions set to zero.
Setup We used 68 subsets in each training iteration, each consisting of two classes. This corresponds to choosing $\rho_{\text{min}}=0.1$ (desired sensitivity, regardless of the true unknown parameter $\rho\in\{0.05,0.1,0.3\}$), with a low $\alpha$ value of 0.5, resulting in the construction of fewer environments according to Equation 10. For each class, we sampled $102r=$ 10 pairs of data points. The representation was defined as $g(z)=wz$ for $w\in\mathbb{R}^{d\times p}$  [^3] A linear representation is chosen to facilitate an analysis of the learned representation space.. Here we focus on the case of $p=16$, $\nu_{z}=\nu_{0}=1$, $\nu^{-}=0.1$, $\nu_{+}=2$, $d_{0}=5$, $d_{1}=d_{2}=10$. The results for additional representation sizes $p$, noise ratios $\frac{\nu^{+}}{\nu^{-}}$ and varying proportions of positive and negative examples are presented in Appendix D.1.
@@
Results In Figure 5 we examine the learned representation. The analysis indicates that ERM prioritizes dimensions 5-15, providing good separation for $a_{1}$, the dominant type in training,
Figure 5: Average feature importance for $\rho=0.9$, 10 repetitions. Our VarAUC penalty favors shared features (blocks 1 and 3), while deprioritizing majority features (block 2). All methods assign low weight to noise features (block 4).
@@
In Figure 4, the learning progress is depicted for $\rho=0.9$ (a similar analysis for $\rho=0.95$ and $\rho=0.7$ can be found in Appendix D). Performance on the same distribution as the training data is similar for ERM and our algorithm, suggesting that applying our algorithm does not negatively impact performance when no distribution shift occurs. However, when there is a distribution shift our algorithm achieves much better results. The VarREx penalty achieves high AUC values more quickly than the VarAUC penalty, but the VarAUC penalty attains higher overall accuracy. IRM shows noisier convergence, since it is applied directly on the gradients, which have been shown to be noisy in contrastive learning due to high variance in data-pair samples [34]. Means and standard deviations are reported in Appendix D.1, as well as the results for additional data dimensions, positive proportions, and variance ratios.
@@ 5.2 Experiments on Real Data
Figure 6: Average percentage changes of our method compared to ERM across 10 repetitions are shown for the ETHEC (top) and CelebA (bottom) datasets. Error bars represent $\pm$ one std-dev.
Experiment 1 - Species Recognition We used the ETHEC dataset [11] which contains 47,978 butterfly images from six families and 561 species (example of the images are provided in Appendix D). We filtered out species with less than five images and focused on images of butterflies from the Lycaenidae and Nymphalidae families. In the training set, $10\%$ of the species were from the Nymphalidae family, while at test time, $90\%$ of the species were from the Nymphalidae family. For each class we sampled $202r=$ 20 pairs.
Experiment 2 - Face Recognition We used the CelebA dataset [30] which contains 202,599 images of 10,177 celebrities. We filtered out people for which the dataset contains less than three images. Following Vinyals et al. [51], we implemented $g$ as a convolutional neural network which has four modules with $3\times 3$ convolutions and 64 filters, followed by batch normalization, a ReLU activation, $2\times 2$ max-pooling, and a fully connected layer of size 32. We used the attribute blond hair for the class distribution shift: for training, we mainly sampled people without blond hair ($95\%$), while at test time, most people ($95\%$) had blond hair. Each training iteration had 150 synthetic environments of two classes and $202r=$ 20 data points per class.
We trained the models on 200 synthetic environments at a time, each of two classes. We implemented $g$ as a fully connected neural network with layers of sizes 128, 64, 32 and 16, and ReLU activations between them.
Experimental Results As can be seen in Figure 6, while all versions of our algorithm show some improvement over ERM, the best results are achieved with the VarAUC penalty (exact means and standard deviations are reported in Table 3 in Appendix D). One-sided paired t-tests show that the improvement over ERM achieved by our algorithm with the VarAUC penalty is statistically significant, with p-values of $<0.04$ on both datasets; p-values for other penalties are reported in Table 4. All p-values were adjusted with FDR [4] correction.
In Appendix D we also provide additional analysis confirming that the main improvement of our algorithm over the ERM baseline stems from improved performance on negative minority pairs.